# Test-time Adaptation for Regression by Subspace Alignment

### v1

@@ 5.2 Test-time Training
Test-time training (TTT) is also similar to TTA as it adapts models with unlabeled target data. The main difference is that TTT requires to modify the model architecture and training procedure in the source domain. The main approach of TTT is additionally training an self-supervised branch simultaneously with the main supervised task in the source domain. Then, during adaptation, the model is updated via minimizing the self-supervised loss on the target data. On the basis of this approach, TTT methods with various self-supervised tasks have been proposed such as rotation prediction [53], contrastive learning [41], clustering [22], or distribution modeling with normalizing flow [48]. However, adding additional losses to the training prohibits the use of off-the-shelf pre-trained models or may potentially affect the performance of the main task. In contrast, TTA accepts arbitrary training methods in the source domain and thus off-the-shelf-models can be adapted.
@@ 5.3 Test-time Adaptation
Test-time adaptation (TTA) aims to adapt a model trained on the source domain to the target domain without accessing the source data [40]. TTA can be applied in a wider range of situations than SFDA and TTT in that TTA does not train additional models or modify the model architecture and source pre-training. TTA for classification has attracted attention for its practicality. Various types of TTA methods have been developed.
Entropy-based. Wang et al. [54] found that the entropy of prediction strongly correlates with accuracy on the target domain and proposed test-time entropy minimization (Tent), which is the most representative of the TTA methods. BACS [61], MEMO [57], EATA [46] and DELTA [60] follow the idea of Tent and improve adaptation performance. T3A [28] adjusts the prototype in the feature space during testing. However, these TTA methods are designed for classification and cannot be applied to regression. For instance, computing entropy, which is widely adopted in TTA, requires a predictive probability for each class, whereas ordinary regression models only output a single predicted value. Thus, we investigate an approach that does not rely on entropy.
Feature alignment. Another approach of TTA is feature alignment. It is based on the insight of UDA and makes the target feature distribution close to the source one. Since accessing the source data is restricted in the TTA setting, methods based on the feature alignment match the statistics of the target features to those of the pre-computed source. BN-adapt [4] updates the feature mean and variance stored in batch normalization (BN) layers [26]. DELTA [60] modifies BN and introduces class-wise loss re-weighting. CFA [31], BUFR [16], CAFe [1], and CAFA [29] incorporate pre-computed source statistics. Although some of these methods are directly applicable to regression, we have observed that they are not effective or even degrade regression performance.
Other tasks. TTA for depth estimation [38] and super resolution [12] are proposed. But they have task-specific architectures or methods and cannot be applied to ordinary regression.
@@ 6 Conclusion
We proposed significant-subspace alignment (SSA), a novel test-time adaptation method for regression models. Since we have found that the naive feature alignment fails in regression TTA because of the learned features being distributed in a small subspace, we incorporated subspace detection and dimension weighting procedures into SSA. Experimental results show that SSA achieved higher $R^{2}$ scores on various regression tasks than did baselines that were originally designed for classification tasks. We will extend TTA to further broader tasks in the future.
@@ Appendix A Limitation
One limitation of SSA is that it assumes a covariate shift, where $p(y|\mathbf{x})$ does not change. Addressing distribution shifts where $p(y|\mathbf{x})$ changes, e.g., concept drift, will be tackled in future work.
@@ Appendix C Experimental Settings
We used PyTorch [49] and PyTorch-Ignite [19] to make the implementations of the source pre-training, proposed method, and baselines. We conducted the experiments with a single NVIDIA A100 GPU.
@@ C.1 Datasets
SVHN [44]: We downloaded SVHN via torchvision.datasets.SVHN. It can be used for non-commercial purposes only [^1] http://ufldl.stanford.edu/housenumbers/.
MNIST [36]: We downloaded MNIST via torchvison.datasets.MNIST. We could not find any license information for MNIST.
UTKFace [59]: We downloaded UTKFace via the official site [^2] https://susanqq.github.io/UTKFace/. It is available for non-commercial research purposes only.
Biwi Kinect [18]: We downloaded Biwi Kinect via Kaggle [^3] https://www.kaggle.com/datasets/kmader/biwi-kinect-head-pose-database. It is released under Attribution-NonCommercial-NoDerivatives 4.0 International (CC BY-NC-ND 4.0) license.
California Housing [47]: We downloaded the dataset from Kaggle [^4] https://www.kaggle.com/datasets/camnugent/california-housing-prices. It is released in public domain.
@@ C.2 Pre-training on the Source Domain
SVHN: We trained ResNet-26 [23] from scratch. For optimization, we used Adam [30] and set the learning rate to 0.0001, weight decay to 0.0005, batch size to 64, and number of epochs to 100. We minimized the MSE loss between the predicted values and digit labels. For the implementation of ResNet-26, we used the PyTorch Image Models (timm) library [55].
UTKFace: We randomly split the dataset into 80% for training and 20% for validation. We fine-tuned ResNet-50 pre-trained on ImageNet [11]. We minimized the MSE loss to predict the ages of the persons in the images. For optimization, we used the same hyperparameters as the above SVHN case. For the implementation of ResNet-50, we used torchvision [42] with IMAGENET1K_V2 initial weights.
@@ C.3 Test-time Adaptation
As for setting the hyperparameters of the baseline methods, we basically followed their original papers. For the evaluation, we ran each TTA three times with different random seeds and reported the means of the scores.
@@
BN-adapt [4]: updates the feature mean and variance stored in the batch normalization layers during testing, i.e., ran inference with model.train() mode in PyTorch.
@@
Feature restoration (FR) [16]: uses the source statistics of the features and outputs as a form of dimension-wise histogram and aligns the target feature histogram to the source one. The original FR uses the histograms of the features and logits pre-computed with the source dataset. Since our focus is on regression models, we used the outputs instead of logits. We set the number of bins of the histograms to eight and the temperature $\tau$ of soft-binning to 0.01 following Eastwood et al. [16]. We set learning rate to 0.0001 (this value gave the best score).
@@
Prototype: We tweaked T3A [28], which regards the weights of the last fully-connected layer as the prototype of each class and updates the prototypes with the mean of the arriving target features during testing. We regarded $\mathbf{w}$ of the linear regressor $h_{\psi}$ as a single prototype and updated it with the mean of the target features. Although T3A determines whether to use a feature for making an update by using entropy, we omitted this component since we cannot compute entropy with regression models.
@@
Variance minimization (VM): makes augmented views of input images and minimizes output variance. This is a modification of MEMO [57], which minimizes the marginal entropy of the augmented views of inputs in the same manner. We set the number of augmented views to 32 per input and set the learning rate to 0.001 with Adam optimizer.
@@
Representation subspace distance (RSD) [6]: The original RSD is a UDA method designed for regression, which aligns the SVD bases between source and target mini-batches. To adapt to TTA, we pre-computed the SVD of the source features wit the whole source dataset instead of mini-batches before TTA and align the target feature SVD with the RSD loss during testing.
@@
SSA (ours): Algorithm 1 lists the procedure of SSA. For optimization, we used Adam [30] with a learning rate$=0.001$, $(\beta_{1},\beta_{2})=(0.9,0.999)$, and weight decay$=0$, which is the default setting in PyTorch [49]. We set the batch size to 64 following other TTA baselines.
@@
Test-time training (TTT) [53]: incorporates a self-supervised rotation prediction task during pre-training in the source domain; then it updates the feature extractor by minimizing the self-supervised loss during testing. The rotation-prediction branch is a linear layer that takes a feature $\mathbf{z}$ and outputs four logits corresponding to the rotation angles $\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$ of an input image. For optimization, we used SGD and set the learning rate to 0.001 in accordance with Sun et al. [53].
@@
Domain adversarial neural network (DANN) [21]: is an unsupervised domain adaptation method which adversarially trains a feature extractor and domain discriminator to learn domain-invariant features. We trained the domain discriminator during training in addition to the main regression model. We used layer4 of ResNet for the discriminator. We scheduled the learning rate and the weight of the domain adaptation loss by following Ganin et al. [21].
@@
  Compute weights for each dimension of the source subspace $\alpha_{d}$ according to Eq. 6
@@
     Project target features $\{\mathbf{z}_{i}^{\text{t}}\}_{i}$ into $\{\tilde{\mathbf{z}}_{i}^{\text{t}}\}_{i}$ according to Eq. 7
     Compute projected target mean ${\tilde{\bm{\mu}}^{\text{t}}}$ and variances $\tilde{\bm{\sigma}}^{\text{t~{}2}}$ analogously to Eqs. 1 and 2
     Update feature extractor $g_{\phi}$ to minimize $\mathcal{L}_{\text{TTA}}$ according to Eq. 8
@@ D.1 Metric for Feature Alignment
We used the KL divergence between two Gaussian distributions in Eq. 4 for the feature alignment in SSA. One may suppose that other metrics could also be used, since the variance term in the denominator makes the naive TTA loss in Eq. 3 unstable, as mentioned in Sec. 3.1. Here, we tried the 2-Wasserstein distance (2WD) between two Gaussian distributions [14] and the L1 norm of the statistics:
@@
We replaced the KL divergence with the 2WD and L1 in Eq. 8 as follows:
$\mathcal{L}_{\text{TTA-2WD}}$ $=\sum_{d=1}^{K}\alpha_{d}W_{2}^{2}\left(\mathcal{N}(\tilde{\mu}_{d}^{\text{t}},\tilde{\sigma}_{d}^{\text{t~{}2}}),\mathcal{N}(0,\lambda_{d}^{\text{s}})\right)$
$=\sum_{d=1}^{K}\alpha_{d}\left\{(\tilde{\mu}_{d}^{\text{t}})^{2}+\!\left(\!\sqrt{\tilde{\sigma}_{d}^{\text{t~{}2}}}-\sqrt{\lambda_{d}^{\text{s}}}\right)^{2}\right\},$ (15)
$\mathcal{L}_{\text{TTA-L1}}$ $=\sum_{d=1}^{K}\alpha_{d}L_{1}\left(\mathcal{N}(\tilde{\mu}_{d}^{\text{t}},\tilde{\sigma}_{d}^{\text{t~{}2}}),\mathcal{N}(0,\lambda_{d}^{\text{s}})\right)$
$=\sum_{d=1}^{K}\alpha_{d}\left\{|\tilde{\mu}_{d}^{\text{t}}|+\left|\sqrt{\tilde{\sigma}_{d}^{\text{t~{}2}}}-\sqrt{\lambda_{d}^{\text{s}}}\right|\right\}.$ (16)
@@
Table 10: $R^{2}$ scores of cases using KL divergence and 2-Wasserstein distance for feature alignment in SSA on UTKFace. The top row is our method.
@@
Table 11: $R^{2}$ scores of cases using KL divergence and 2-Wasserstein distance for feature alignment in SSA on Biwi Kinect. The top row is our method.
@@
Tabs. 9, 10 and 11 compare the effects of using the KL divergence and 2WD on SVHN-MNIST, UTKFace, and Biwi Kinect. The KL divergence with subspace detection (SSA) achieved highest $R^{2}$ scores in almost all cases. In contrast, the 2WD variant of SSA produced only a slight improvement over Source on SVHN-MNIST (Tab. 9) and sometimes it had even worse scores than Source on Biwi Kinect (Tab. 11). This degradation of 2WD is because the scale of the variance $\sigma_{d}^{2}$ is different among feature dimensions $d$. The KL divergence can absorb the difference in scale since it includes the ratio of the variances, as in Eq. 8.
@@ D.2 Feature Visualization
Fig. 4 illustrates PCA visualizations of the source and target features after TTA. In the visualizations of SVHN-MNIST (a) and Biwi Kinect (c), we can see that SSA makes the target feature distribution fit within the source distribution while the target features of the baselines protrude from the source distribution. In UTKFace (b), the target features of Source significantly degenerate to a single point, but the other methods alleviate this. In Fig. 4, we also report the optimal transport distance (OTD) between the features of the both domains, which evaluates the feature alignment quantitatively. In addition to that our SSA alleviates the feature distribution gap in the subspace, SSA retains the source subspace better, as shown in Fig. 2 of Sec. 4.3.4.

### v2

@@ 5.2 Test-time Training
Test-time training (TTT) is also similar to TTA as it adapts models with unlabeled target data. The main difference is that TTT requires to modify the model architecture and training procedure in the source domain. The main approach of TTT is additionally training an self-supervised branch simultaneously with the main supervised task in the source domain. Then, during adaptation, the model is updated via minimizing the self-supervised loss on the target data. On the basis of this approach, TTT methods with various self-supervised tasks have been proposed such as rotation prediction (Sun et al., 2020), contrastive learning (Liu et al., 2021), clustering (Hakim et al., 2023), or distribution modeling with normalizing flow (Osowiechi et al., 2023). However, adding additional losses to the training prohibits the use of off-the-shelf pre-trained models or may potentially affect the performance of the main task. In contrast, TTA accepts arbitrary training methods in the source domain and thus off-the-shelf-models can be adapted.
@@ 5.3 Test-time Adaptation
Test-time adaptation (TTA) aims to adapt a model trained on the source domain to the target domain without accessing the source data (Liang et al., 2023). TTA can be applied in a wider range of situations than SFDA and TTT in that TTA does not train additional models or modify the model architecture and source pre-training. TTA for classification has attracted attention for its practicality. Various types of TTA methods have been developed.
Entropy-based. Wang et al. (2021) found that the entropy of prediction strongly correlates with accuracy on the target domain and proposed test-time entropy minimization (Tent), which is the most representative of the TTA methods. BACS (Zhou & Levine, 2021), MEMO (Zhang et al., 2022), EATA (Niu et al., 2022) and DELTA (Zhao et al., 2023) follow the idea of Tent and improve adaptation performance. T3A (Iwasawa & Matsuo, 2021) adjusts the prototype in the feature space during testing. IST (Ma, 2024) employs graph-based pseudo label modification. However, these TTA methods are designed for classification and cannot be applied to regression. For instance, computing entropy, which is widely adopted in TTA, requires a predictive probability for each class, whereas ordinary regression models only output a single predicted value. Thus, we investigate an approach that does not rely on entropy.
Feature alignment. Feature alignment is based on the insight of UDA and makes the target feature distribution close to the source one. Since accessing the source data is restricted in the TTA setting, methods based on the feature alignment match the statistics of the target features to those of the pre-computed source. BN-adapt (Benz et al., 2021) updates the feature mean and variance stored in batch normalization (BN) layers (Ioffe & Szegedy, 2015). DELTA (Zhao et al., 2023) modifies BN and introduces class-wise loss re-weighting. CFA (Kojima et al., 2022), BUFR (Eastwood et al., 2022), CAFe (Adachi et al., 2023), and CAFA (Jung et al., 2023) incorporate pre-computed source statistics. Although some of these methods are directly applicable to regression, we have observed that they are not effective or even degrade regression performance.
Other tasks. TTA for depth estimation (Li et al., 2023) super resolution (Deng et al., 2024), point cloud (Wang et al., 2024), and person re-identification (Adachi et al., 2024) are proposed. But they have task-specific architectures or methods and cannot be applied to ordinary regression.
@@ 6 Conclusion
We proposed significant-subspace alignment (SSA), a novel test-time adaptation method for regression models. Since we have found that the naive feature alignment fails in regression TTA because of the learned features being distributed in a small subspace, we incorporated subspace detection and dimension weighting procedures into SSA. Experimental results show that SSA achieved higher $R^{2}$ scores on various regression tasks than did baselines that were originally designed for classification tasks. We will extend TTA to further broader tasks and settings such as concept drift in the future.
Ethics statement. The potential ethical concern is that a model may have fairness or bias issues in certain sensitive applications if the model adapts to biased target data. The model’s behavior should be carefully monitored in such a situation.
Reproducibility statement. Details on the datasets and experimental settings are described in Sections 4.1, 4.2 and C. We also provide the code in the supplementary material.
@@ Appendix A Limitation
One limitation of SSA is that it assumes a covariate shift, where $p(y|\mathbf{x})$ does not change. Addressing distribution shifts where $p(y|\mathbf{x})$ changes, e.g., concept drift, will specifically be addressed as a target in future work.
@@ Appendix C Experimental Settings
We used PyTorch (Paszke et al., 2019) and PyTorch-Ignite (Fomin et al., 2020) to make the implementations of the source pre-training, proposed method, and baselines. We conducted the experiments with a single NVIDIA A100 GPU.
@@ C.1 Datasets
SVHN (Netzer et al., 2011): We downloaded SVHN via torchvision.datasets.SVHN. It can be used for non-commercial purposes only [^1] http://ufldl.stanford.edu/housenumbers/.
MNIST (LeCun et al., 1998b): We downloaded MNIST via torchvison.datasets.MNIST. We could not find any license information for MNIST.
UTKFace (Zhang et al., 2017): We downloaded UTKFace via the official site [^2] https://susanqq.github.io/UTKFace/. It is available for non-commercial research purposes only.
Biwi Kinect (Fanelli et al., 2013): We downloaded Biwi Kinect via Kaggle [^3] https://www.kaggle.com/datasets/kmader/biwi-kinect-head-pose-database. It is released under Attribution-NonCommercial-NoDerivatives 4.0 International (CC BY-NC-ND 4.0) license.
California Housing (Nugent, 2017): We downloaded the dataset from Kaggle [^4] https://www.kaggle.com/datasets/camnugent/california-housing-prices. It is released in public domain.
@@ C.2 Pre-training on the Source Domain
SVHN: We trained ResNet-26 (He et al., 2016) from scratch. For optimization, we used Adam (Kingma & Ba, 2015) and set the learning rate to 0.0001, weight decay to 0.0005, batch size to 64, and number of epochs to 100. We minimized the MSE loss between the predicted values and digit labels. For the implementation of ResNet-26, we used the PyTorch Image Models (timm) library (Wightman, 2019).
UTKFace: We randomly split the dataset into 80% for training and 20% for validation. We fine-tuned ResNet-50 pre-trained on ImageNet (Deng et al., 2009). We minimized the MSE loss to predict the ages of the persons in the images. For optimization, we used the same hyperparameters as the above SVHN case. For the implementation of ResNet-50, we used torchvision (maintainers & contributors, 2016) with IMAGENET1K_V2 initial weights.
@@ C.3 Test-time Adaptation
As for setting the hyperparameters of the baseline methods, we basically followed their original papers. For adaptation, we adopted an offline manner for fair comparison, i.e., we ran TTA for one epoch and then ran evaluation, which is widely adopted in existing TTA works (Wang et al., 2021; Zhou & Levine, 2021; Eastwood et al., 2022). For the evaluation, we ran each TTA three times with different random seeds and reported the means of the scores.
@@
BN-adapt (Benz et al., 2021): updates the feature mean and variance stored in the batch normalization layers during testing, i.e., ran inference with model.train() mode in PyTorch.
@@
Feature restoration (FR) (Eastwood et al., 2022): uses the source statistics of the features and outputs as a form of dimension-wise histogram and aligns the target feature histogram to the source one. The original FR uses the histograms of the features and logits pre-computed with the source dataset. Since our focus is on regression models, we used the outputs instead of logits. We set the number of bins of the histograms to eight and the temperature $\tau$ of soft-binning to 0.01 following Eastwood et al. (2022). We set learning rate to 0.0001 (this value gave the best score).
@@
Prototype: We tweaked T3A (Iwasawa & Matsuo, 2021), which regards the weights of the last fully-connected layer as the prototype of each class and updates the prototypes with the mean of the arriving target features during testing. We regarded $\mathbf{w}$ of the linear regressor $h_{\psi}$ as a single prototype and updated it with the mean of the target features. Although T3A determines whether to use a feature for making an update by using entropy, we omitted this component since we cannot compute entropy with regression models.
@@
Variance minimization (VM): makes augmented views of input images and minimizes output variance. This is a modification of MEMO (Zhang et al., 2022), which minimizes the marginal entropy of the augmented views of inputs in the same manner. We set the number of augmented views to 32 per input and set the learning rate to 0.001 with Adam optimizer.
@@
Representation subspace distance (RSD) (Chen et al., 2021): The original RSD is a UDA method designed for regression, which aligns the SVD bases between source and target mini-batches. To adapt to TTA, we pre-computed the SVD of the source features wit the whole source dataset instead of mini-batches before TTA and align the target feature SVD with the RSD loss during testing.
@@
SSA (ours): Algorithm 1 lists the procedure of SSA. For optimization, we used Adam (Kingma & Ba, 2015) with a learning rate$=0.001$, $(\beta_{1},\beta_{2})=(0.9,0.999)$, and weight decay$=0$, which is the default setting in PyTorch (Paszke et al., 2019). We set the batch size to 64 following other TTA baselines.
@@
Test-time training (TTT) (Sun et al., 2020): incorporates a self-supervised rotation prediction task during pre-training in the source domain; then it updates the feature extractor by minimizing the self-supervised loss during testing. The rotation-prediction branch is a linear layer that takes a feature $\mathbf{z}$ and outputs four logits corresponding to the rotation angles $\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$ of an input image. For optimization, we used SGD and set the learning rate to 0.001 in accordance with Sun et al. (2020).
@@
Domain adversarial neural network (DANN) (Ganin et al., 2016): is an unsupervised domain adaptation method which adversarially trains a feature extractor and domain discriminator to learn domain-invariant features. We trained the domain discriminator during training in addition to the main regression model. We used layer4 of ResNet for the discriminator. We scheduled the learning rate and the weight of the domain adaptation loss by following Ganin et al. (2016).
@@
  Compute weights for each dimension of the source subspace $\alpha_{d}$ according to Equation 5
@@
     Project target features $\{\mathbf{z}_{i}^{\text{t}}\}_{i}$ into $\{\tilde{\mathbf{z}}_{i}^{\text{t}}\}_{i}$ according to Equation 6
     Compute projected target mean ${\tilde{\bm{\mu}}^{\text{t}}}$ and variances $\tilde{\bm{\sigma}}^{\text{t~{}2}}$ analogously to Equation 1
     Update $\phi$ of the feature extractor $g_{\phi}$ to minimize $\mathcal{L}_{\text{TTA}}$ according to Equation 7
@@ D.1 Metric for Feature Alignment
We used the KL divergence between two Gaussian distributions in Equation 3 for the feature alignment in SSA. One may suppose that other metrics could also be used, since the variance term in the denominator makes the naive TTA loss in Equation 2 unstable, as mentioned in Section 3.1. Here, we tried the 2-Wasserstein distance (2WD) between two Gaussian distributions (Dowson & Landau, 1982) and the L1 norm of the statistics:
@@
We replaced the KL divergence with the 2WD and L1 in Equation 7 as follows:
$\mathcal{L}_{\text{TTA-2WD}}\!=\!\sum_{d=1}^{K}\!\alpha_{d}W_{2}^{2}\!\left(\mathcal{N}(\tilde{\mu}_{d}^{\text{t}},\tilde{\sigma}_{d}^{\text{t~{}2}}),\mathcal{N}(0,\lambda_{d}^{\text{s}})\right)\!$ $=\!\!\sum_{d=1}^{K}\!\alpha_{d}\!\left\{\!(\tilde{\mu}_{d}^{\text{t}})^{2}\!+\!\left(\!\sqrt{\tilde{\sigma}_{d}^{\text{t~{}2}}}-\!\sqrt{\lambda_{d}^{\text{s}}}\right)^{\!2}\right\},$ (13)
$\mathcal{L}_{\text{TTA-L1}}\!=\!\sum_{d=1}^{K}\!\alpha_{d}L_{1}\left(\mathcal{N}(\tilde{\mu}_{d}^{\text{t}},\tilde{\sigma}_{d}^{\text{t~{}2}}),\mathcal{N}(0,\lambda_{d}^{\text{s}})\right)\!$ $=\!\!\sum_{d=1}^{K}\!\alpha_{d}\!\left\{|\tilde{\mu}_{d}^{\text{t}}|+\left|\sqrt{\tilde{\sigma}_{d}^{\text{t~{}2}}}-\sqrt{\lambda_{d}^{\text{s}}}\right|\right\}.$ (14)
@@
Table 10: Test $R^{2}$ scores of cases using KL divergence and 2-Wasserstein distance for feature alignment in SSA on UTKFace. The top row is our method.
@@
Table 11: Test $R^{2}$ scores of cases using KL divergence and 2-Wasserstein distance for feature alignment in SSA on Biwi Kinect. The top row is our method.
@@
Tables 9, 10 and 11 compare the effects of using the KL divergence and 2WD on SVHN-MNIST, UTKFace, and Biwi Kinect. The KL divergence with subspace detection (SSA) achieved highest $R^{2}$ scores in almost all cases. In contrast, the 2WD variant of SSA produced only a slight improvement over Source on SVHN-MNIST (Table 9) and sometimes it had even worse scores than Source on Biwi Kinect (Table 11). This degradation of 2WD is because the scale of the variance $\sigma_{d}^{2}$ is different among feature dimensions $d$. The KL divergence can absorb the difference in scale since it includes the ratio of the variances, as in Equation 7.
@@ D.2 Feature Visualization
Source Refer to caption Refer to caption Refer to caption
SSA (Ours) Refer to caption Refer to caption Refer to caption
(a) SVHN-MNIST (b) UTKFace (c) Biwi Kinect
Figure 5: UMAP (McInnes et al., 2018) visualizations of source and target features on each dataset. The blue and orange dots represent the source and target features, respectively.
Figure 4 illustrates PCA visualizations of the source and target features after TTA. In the visualizations of SVHN-MNIST (a) and Biwi Kinect (c), we can see that SSA makes the target feature distribution fit within the source distribution while the target features of the baselines protrude from the source distribution. In UTKFace (b), the target features of Source significantly degenerate to a single point, but the other methods alleviate this. In Figure 4, we also report the optimal transport distance (OTD) between the features of the both domains, which evaluates the feature alignment quantitatively. In addition to that our SSA alleviates the feature distribution gap in the subspace, SSA retains the source subspace better, as shown in Figure 2 of Section 4.3.4.
We also visualized the source and target features with UMAP (McInnes et al., 2018) in Figure 5 to see the relation not limited within the top-2 principal components. We first trained the UMAP mapping with the source features projected onto the top $K=100$ dimensional principal component space before TTA, which are shared among TTA methods. Then, we mapped the target features after TTA with the learned UMAP mapping. We can also see that the target feature distribution becomes closer to the source one by our SSA.